OpenStudy (he66666):

Grade 12 math: exponential & logarithmic function? Dry cleaners use a cleaning fluid that is purified by evaporation and condensation after each cleaning cycle. Every time the fluid is purified, 2.1% of it is lost. The fluid has to be topped up when half of the original fluid remains. After how many cycles will the fluid need to be topped up? Answer: 33 cycles How do you solve this question?

OpenStudy (anonymous):

losses 2.1% means multiply by .979

OpenStudy (anonymous):

it is like a compound interest problem, but different. sorry, but i dont really know

OpenStudy (anonymous):

each cycle so solve \[.979^x =\frac{1}{2}\]

OpenStudy (anonymous):

you have a choice. you can use logs, or you can just guess and check but obviously you are supposed to use logs so lets do it

OpenStudy (anonymous):

\[.979^x=.5\] \[x\ln(.979)=\ln(.5)\] \[x=\frac{\ln(.5)}{ln(.979)}\]

OpenStudy (anonymous):

close to 33

OpenStudy (he66666):

But why is it .979? Shouldn't it be 99.979 since you do 100-0.021? :S

OpenStudy (anonymous):

hope it is clear that if you lose 2.1% you retain 97.9 percent

OpenStudy (anonymous):

oh wait. in math we use numbers not their percent form to compute with. 100% =1, 2.1% =.021

OpenStudy (anonymous):

you should subtract .021 from 1 or subtract 2.1 from 100 and then convert to a decimal

OpenStudy (he66666):

Oh okay. Thanks so much! :)

OpenStudy (anonymous):

welcome